# Panning for Gold: Expanding Domain-Specific Knowledge Graphs with General Knowledge

### v1

@@ 3. Problem statement
3.1. Domain-specific Knowledge Graph
The domain-specific knowledge graphs $G^{d}=(V^{d},R^{d},F^{d})$ and general knowledge graphs $G^{g}=(V^{g},R^{g},F^{g})$ encapsulate real-world domain-specific and general knowledge through three core components: entity sets $V^{d},V^{g}$, relation sets $R^{d},R^{g}$, and fact sets $F^{d},F^{g}$ (Arenas et al., 2021). Each fact in $F^{d}$ and $F^{g}$ is represented in the form of $(v^{d},r^{d},\hat{v}^{d})\in V^{d}\times R^{d}\times V^{d}$ and $(v^{g},r^{g},\hat{v}^{g})\in V^{g}\times R^{g}\times V^{g}$, respectively, where entities $v^{d},\hat{v}^{d}\in V^{d}$, $v^{g},\hat{v}^{g}\in V^{g}$, relations $r^{d}\in R^{d}$, and $r^{g}\in R^{g}$.
3.2. Domain-specific Knowledge Graph Fusion
Based on the aforementioned definitions, the domain-specific knowledge graph fusion task aims to mine the highly relevant knowledge in the general knowledge graph $G^{g}=\big(V^{g},R^{g},F^{g}\big)$ to enhance the knowledge coverage of domain-specific knowledge graph $G^{d}=\big(V^{d},R^{d},F^{d}\big)$. Formally, given a DKG $G^{d}$ and GKG $G^{g}$, the DKGF task generates a new set of fused facts $F_{fused}$ for $G^{d}$:
@@ 4. Methodology
In this section, we present the ExeFuse framework, which introduces a novel Fact-as-Program paradigm for domain-specific knowledge graph fusion. We first introduce the core insight, then present the unified probabilistic framework, and finally describe the model architecture in detail.
4.1. Core Insight: Relevance as Executability
The fundamental insight underlying our approach is that domain relevance can be operationalized as program executability. Specifically, we propose that a general fact $f^{g}=(v^{g},r^{g},\hat{v}^{g})$ is domain-relevant to DKG $G^{d}$ if and only if there exists a semantic operator $z$ such that the semantic program $\pi(f^{g},z)$ induced by $f^{g}$ and $z$ can be executed on $G^{d}$ to produce valid domain facts.
This insight addresses Challenge I by providing a principled, executable-based criterion for domain relevance assessment. Instead of relying on semantic similarity measures that may fail when explicit connections are absent (e.g., “Brookings Institution” and “Iran” have low semantic similarity but are highly relevant in political crisis analysis), we verify relevance by checking whether the induced program can generate valid domain facts through execution on the DKG structure.
4.2. Unified Probabilistic Framework
To operationalize the Fact-as-Program paradigm, we formulate the DKGF task within a unified probabilistic framework. Given a general fact $f^{g}=(v^{g},r^{g},\hat{v}^{g})$ and DKG $G^{d}$, we aim to estimate the probability of generating a domain fact $f^{d}=(v^{d},r^{d},\hat{v}^{d})$ that should be fused into $G^{d}$:
(2) $p(f^{d}\mid f^{g},G^{d})=\sum_{z\in\mathcal{Z}}p(z\mid f^{g})\cdot p(f^{d}\mid z,f^{g},G^{d}),$
where $\mathcal{Z}$ is the set of semantic operators. The framework naturally decomposes the two challenges: $p(z\mid f^{g})$ addresses Challenge II by mapping abstract GKG relations to executable semantic operators at appropriate granularity levels, while $p(f^{d}\mid z,f^{g},G^{d})$ addresses Challenge I by verifying executability and generating domain-relevant facts.
4.3. Model Architecture
Figure 2 illustrates the overall architecture of ExeFuse. The framework consists of three main stages: ① operator mapping and program composition, which maps GKG relations to semantic operators and composes them into executable programs; ② program execution, which executes the composed program on the DKG structure; and ③ executability verification, which generates domain facts and verifies program executability to assess domain relevance.
4.3.1. Stage 1: Operator Mapping and Program Composition
Given a GKG relation $r^{g}$, we first embed it into a $d$-dimensional space: $\mathbf{r}_{e}=\text{Emb}(r^{g})\in\mathbb{R}^{d}$, where $\text{Emb}(\cdot)$ is a learnable relation embedding matrix.
Granularity-Aware Operator Mapping.  To address Challenge II (granularity misalignment), we map $r^{g}$ to semantic operators at three granularity levels (coarse, medium, fine). This allows the model to handle the abstraction gap between GKG and DKG relations. For example, a coarse-grained GKG relation like “head of government” can be mapped to fine-grained operators that generate specific actions like ”make_optimistic_comment”.
First, we compute the granularity level distribution:
(3) $\mathbf{g}=\text{Softmax}(\mathbf{W}_{g}\mathbf{r}_{e}+\mathbf{b}_{g})\in\mathbb{R}^{3},$
where $\mathbf{W}_{g}\in\mathbb{R}^{3\times d}$ and $\mathbf{b}_{g}\in\mathbb{R}^{3}$ are learnable parameters, and $\mathbf{g}=[g_{\text{coarse}},g_{\text{medium}},g_{\text{fine}}]^{T}$ represents the probability distribution over the three granularity levels.
For each granularity level $l\in\{\text{coarse},\text{medium},\text{fine}\}$, we maintain a set of operators $\mathcal{Z}_{l}$ with $|\mathcal{Z}_{l}|=n_{l}$ operators. The operator selection at level $l$ is computed as:
(4) $\mathbf{w}_{l}=\text{Softmax}(\mathbf{W}_{l}\mathbf{r}_{e}+\mathbf{b}_{l})\in\mathbb{R}^{n_{l}},$
where $\mathbf{W}_{l}\in\mathbb{R}^{n_{l}\times d}$ and $\mathbf{b}_{l}\in\mathbb{R}^{n_{l}}$ are level-specific learnable parameters.
The semantic operator embedding is then computed as a weighted combination:
(5) $\mathbf{z}_{\text{sem}}=\sum_{l\in\{\text{coarse},\text{medium},\text{fine}\}}g_{l}\cdot\sum_{i=1}^{n_{l}}w_{l,i}\cdot\mathbf{z}_{l,i}\in\mathbb{R}^{d},$
where $\mathbf{z}_{l,i}\in\mathbb{R}^{d}$ is the embedding of the $i$-th operator at granularity level $l$, and $w_{l,i}$ is the $i$-th element of $\mathbf{w}_{l}$.
Relevance Operator Mapping.  In parallel, we map $r^{g}$ to domain relevance operators, which are specifically designed to assess domain relevance: $\mathbf{z}_{\text{rel}}=\sum_{i=1}^{n_{\text{rel}}}w_{\text{rel},i}\cdot\mathbf{z}_{\text{rel},i}$, where $\mathbf{w}_{\text{rel}}=\text{Softmax}(\mathbf{W}_{\text{rel}}\mathbf{r}_{e}+\mathbf{b}_{\text{rel}})$ and $n_{\text{rel}}$ is the number of relevance operators.
Program Composition.  The semantic and relevance operators are composed using multi-head attention to form a unified program representation $\pi_{\text{final}}\in\mathbb{R}^{d}$. This composition enables the model to leverage both granularity-aware semantic information and domain relevance information.
Granularity Transformation.  We transform the relation embedding based on the selected granularity level using a Transformer encoder: $\mathbf{r}_{\text{transformed}}=\text{TransformerEncoder}(\mathbf{r}_{e}+\mathbf{e}_{l})$, where $\mathbf{e}_{l}\in\mathbb{R}^{d}$ is a learnable granularity embedding for level $l$, and $l^{*}=\arg\max_{l}g_{l}$ is the selected granularity level.
4.3.2. Stage 2: Program Execution
The composed program is executed on the DKG structure through program-relation interaction. Given the transformed relation embedding $\mathbf{r}_{\text{transformed}}$ and program embedding $\pi_{\text{final}}$, we compute the executed relation representation:
(6) $\mathbf{r}_{\text{exec}}=\text{MLP}([\mathbf{r}_{\text{transformed}};\pi_{\text{final}}])+\mathbf{r}_{\text{transformed}}\in\mathbb{R}^{d},$
where $\cdot$$\cdot$] denotes concatenation, MLP is a multi-layer perceptron, and the residual connection preserves the original relation information. This execution step effectively “runs” the semantic program on the DKG structure, producing a relation representation that is adapted to the domain context.
4.3.3. Stage 3: Executability Verification
The core innovation of our approach is the Domain Fact Generator, which directly implements the insight that relevance equals executability. Given the program embedding $\pi_{\text{final}}$, head entity $v^{g}$, tail entity $\hat{v}^{g}$, and DKG $G^{d}$, the generator produces a probability distribution over domain relations $R^{d}$:
(7) $p(r^{d}\mid z,v^{g},\hat{v}^{g},G^{d})=\text{Softmax}(\text{MLP}([\pi_{\text{final}};\mathbf{v}_{e}^{g};\hat{\mathbf{v}}_{e}^{g}]))\in\mathbb{R}^{|R^{d}|},$
where $\mathbf{v}_{e}^{g}=\text{Emb}(v^{g})$ and $\hat{\mathbf{v}}_{e}^{g}=\text{Emb}(\hat{v}^{g})$ are entity embeddings. The generator network consists of multiple MLP layers with layer normalization and GELU activation.
The executability score is computed as:
(8) $s_{\text{exec}}=p(r^{g}\mid z,v^{g},\hat{v}^{g},G^{d}),$
which measures the probability that the generator assigns to the original GKG relation $r^{g}$ in the generated distribution over $R^{d}$. A high score indicates that the program can generate the original relation, thus verifying domain relevance. This score directly operationalizes our core insight: if a program is executable (i.e., can generate valid domain facts), the corresponding GKG fact is domain-relevant.
4.3.4. Final Prediction
The final fusion prediction combines the executability score with structural consistency. We compute a space distance to measure structural consistency in the embedding space:
(9) $d_{\text{space}}=\|\mathbf{v}_{e}^{g}+\mathbf{r}_{\text{exec}}-\hat{\mathbf{v}}_{e}^{g}\|_{1},$
which is then converted to a similarity score: $s_{\text{struct}}=\exp(-d_{\text{space}}/\tau)$, where $\tau>0$ is a temperature parameter. The final prediction is:
@@ 4.3.4. Final Prediction
(10) $p(\text{fuse}\mid f^{g},G^{d})=\text{Sigmoid}(\text{MLP}([\mathbf{v}_{e}^{g};\hat{\mathbf{v}}_{e}^{g};\mathbf{r}_{\text{exec}};s_{\text{exec}};s_{\text{struct}}]))\in[0,1],$
where the MLP fuses all signals: entity embeddings, executed relation representation, executability score, and structural consistency score. The executability score captures semantic relevance (addressing Challenge I), while the structural score captures embedding-space consistency.
@@ 4.4. Training Objective
Given a training set $\mathcal{D}=\{(f_{i}^{g},y_{i})\}_{i=1}^{N}$ where $y_{i}\in\{0,1\}$ indicates whether $f_{i}^{g}$ should be fused, we minimize:
(11) $\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right]+\beta\mathcal{L}_{\text{aux}},$
where $p_{i}=p(\text{fuse}\mid f_{i}^{g},G^{d})$ is the predicted fusion probability, and $\mathcal{L}_{\text{aux}}$ includes consistency and contrastive losses to encourage alignment between executability and structural scores. Optionally, we use Focal Loss to address class imbalance.
Table 1. Dataset statistics. “#Ent”, “#Rel.”: The number of entities and relations in DKG and GKG, respectively. “Structure. Sim.”: The average neighbor structure similarity of entities in DKG and GKG, as defined in (Jiang et al., 2024c). “#Rel. O.”: The proportion of overlapping relations in DKG and GKG, respectively.

### v2

@@ 3. Problem Statement
Domain-specific Knowledge Graph.  The domain-specific knowledge graphs $G^{d}=(V^{d},R^{d},F^{d})$ and general knowledge graphs $G^{g}=(V^{g},R^{g},F^{g})$ encapsulate real-world domain-specific and general knowledge through three core components: entity sets $V^{d},V^{g}$, relation sets $R^{d},R^{g}$, and fact sets $F^{d},F^{g}$ (Wang et al., 2025). Each fact in $F^{d}$ and $F^{g}$ is represented in the form of $(v^{d},r^{d},\hat{v}^{d})\in V^{d}\times R^{d}\times V^{d}$ and $(v^{g},r^{g},\hat{v}^{g})\in V^{g}\times R^{g}\times V^{g}$, respectively, where entities $v^{d},\hat{v}^{d}\in V^{d}$, $v^{g},\hat{v}^{g}\in V^{g}$, relations $r^{d}\in R^{d}$, and $r^{g}\in R^{g}$.
Domain-specific Knowledge Graph Fusion.  Based on the aforementioned definitions, the domain-specific knowledge graph fusion task aims to mine the highly relevant knowledge in the general knowledge graph $G^{g}=\big(V^{g},R^{g},F^{g}\big)$ to enhance the knowledge coverage of domain-specific knowledge graph $G^{d}=\big(V^{d},R^{d},F^{d}\big)$. Formally, given a DKG $G^{d}$ and GKG $G^{g}$, the DKGF task generates a new set of fused facts $F_{fused}$ for $G^{d}$:
@@ 4. Methodology
In this section, we present ExeFuse, a neuro-symbolic framework designed to address the unique challenges of domain-specific knowledge graph fusion (DKGF). As analyzed in Section 1, the core difficulty of DKGF lies in high ambiguity of domain relevance and cross-domain granularity misalignment. Existing static similarity matching fails to bridge this gap. Instead, effective fusion requires answering two fundamental questions: (i) how to uncover implicit logical connections hidden beneath surface-level dissimilarity, and (ii) how to adapt abstract general facts into the specific granularity of the target domain.
To this end, we propose the Fact-as-Program paradigm, which reframes fusion not as a classification task, but as a dynamic execution process (Pojer et al., 2025; Cheng et al., 2024; Alshahrani et al., 2017; Liu et al., 2025; Cheng and Sun, 2024). As illustrated in Figure 2, ExeFuse consists of three coordinated stages:
Stage 1: Latent Program Induction.  To handle structural heterogeneity, we first encode discrete GKG facts into continuous program states using a structure-aware predicate, preparing them for logical transformation.
Stage 2: Resolving Ambiguity via Neuro-Symbolic Execution.  To resolve the ambiguity of relevance, we move beyond surface embedding similarity. We treat logic rules as transition operators that “execute” the fact. By applying these operators, we infer the logical reachability of a fact—transforming it from its original state to a potential domain-relevant state. This ensures that even semantically distant entities can be linked if there exists a valid logical path.
Stage 3: Granularity Calibration via Target Space Grounding:.  To address granularity misalignment, we view the execution result as a proposal that must be verified. We construct a Target State Space representing the valid semantic distribution of the DKG. The “program” is considered successfully executed only if its output can be grounded (i.e., mapped with high probability) into this specific target region, ensuring the fused knowledge aligns with the domain’s granularity.
4.1. Formalization & Overview
In this paradigm, we treat each source GKG fact $f^{g}=(v^{g},r^{g},\hat{v}^{g})$ as a latent executable program $\Pi(f^{g})$ defined in a continuous Quasi-Symbolic Space $\mathcal{Q}\subseteq\mathbb{R}^{d}$. formally, the framework is a tuple $\Pi=(\phi_{\text{enc}},\mathcal{L},\mathcal{Q})$, where:
  • $\phi_{\text{enc}}$ is the Structure-Aware Neural Predicate that maps discrete GKG facts into continuous program states, preserving structural semantics.
  • $\mathcal{L}$ is the set of Inductive Logic Rules acting as transition operators to transform states into domain states.
  • $\mathcal{Q}$ is the differentiable state space where geometric transformations correspond to logical entailments.
A GKG fact $f^{g}$ is deemed valid for fusion into the DKG $G^{d}$ if its program is executable. Executability is defined as the ability to transition from an initial state $\mathbf{q}^{g}$ to a valid target state $\mathbf{q}^{d}$ that resides within the Target State Space of the DKG.
4.2. Latent Program Induction via Structure-Aware Predicates
The first stage compiles the discrete symbolic fact into an initial executable state. Unlike standard approaches that treat facts as flat feature vectors, we propose a Structure-Aware Neural Predicate to capture the compositional logic of the triplet $(v^{g},r^{g},\hat{v}^{g})$.
4.2.1. Structure-Aware Neural Predicate
Let $\mathbf{E}^{g}\in\mathbb{R}^{|V^{g}|\times d_{emb}}$ and $\mathbf{R}^{g}\in\mathbb{R}^{|R^{g}|\times d_{emb}}$ be the pre-trained embedding matrices. We define the initial program state $\mathbf{q}^{g}\in\mathcal{Q}$ via a relational composition function:
$\mathbf{q}^{g}=\phi_{\text{enc}}(f^{g})=\mathbf{W}_{2}\cdot\text{GeLU}\left(\mathbf{W}_{1}\left[(\mathbf{e}_{v^{g}}\odot\mathbf{e}_{r^{g}})\,;\,(\mathbf{e}_{\hat{v}^{g}}\odot\mathbf{e}_{r^{g}})\,;\,\mathbf{e}_{r^{g}}\right]+\mathbf{b}_{1}\right),$
where $\mathbf{e}_{(\cdot)}$ denotes the embedding vector retrieved from $\mathbf{E}^{g}$ or $\mathbf{R}^{g}$, $\odot$ represents the Hadamard (element-wise) product, and $[;]$ denotes concatenation. $\mathbf{W}_{1},\mathbf{W}_{2}$ and $\mathbf{b}_{1}$ are learnable parameters. Remark: The interaction terms $(\mathbf{e}_{v}\odot\mathbf{e}_{r})$ and $(\mathbf{e}_{\hat{v}}\odot\mathbf{e}_{r})$ explicitly model the interaction between the relation and the head/tail entities, ensuring that $\mathbf{q}^{g}$ encodes the structural validity of the fact.
4.2.2. Target State Space
We define the Target State Space $\mathcal{Q}^{d}\subset\mathcal{Q}$ as the semantic region populated by the embeddings of all existing valid DKG facts:
(3) $\mathcal{Q}^{d}=\{\phi_{\text{dkg}}(f^{d})\mid f^{d}\in F^{d}\},$
where $\phi_{\text{dkg}}$ is the domain-specific encoder. It shares the same architecture as $\phi_{\text{enc}}$ but is fine-tuned to represent the target domain distribution.
4.3. Stage 2: Resolving Ambiguity via Neuro-Symbolic Execution
This stage addresses Challenge I (Ambiguity of Domain Relevance). In a static space, relevance is often ambiguous because direct similarity fails to capture implicit connections. We argue that relevance is best established through logical reachability (Spillo et al., 2024): if a fact implies a valid domain state through explicit reasoning rules, the ambiguity is resolved.
4.3.1. Differentiable Logic Rules
We learn a set of logic rules $\mathcal{L}$. Symbolically, a rule $R\in\mathcal{L}$ implies $r^{d}(X,Y)\leftarrow r^{g}(X,Y)\wedge\mathcal{C}$. To support differentiable execution, we parameterize each rule $R$ as an affine transformation. The rule application function $\mathcal{T}_{R}:\mathcal{Q}\to\mathcal{Q}$ is defined as:
(4) $\mathcal{T}_{R}(\mathbf{q}^{g})=\mathbf{w}_{R}\odot\mathbf{q}^{g}+\mathbf{b}_{R},$
where $\mathbf{w}_{R},\mathbf{b}_{R}\in\mathbb{R}^{d}$ are learnable parameters. Geometrically, this transformation models the inference step, where $\mathbf{w}_{R}$ and $\mathbf{b}_{R}$ adjust the semantic focus to trace the logical path from the source fact to the target context.
4.3.2. Neural Rule Selector
To address the ambiguity of which rule applies, we employ a Neural Rule Selector that computes an attention distribution $p(R\mid\mathbf{q}^{g})$ over $\mathcal{L}$:
(5) $p(R\mid\mathbf{q}^{g})=\frac{\exp(\mathbf{u}_{R}^{\top}\tanh(\mathbf{W}_{sel}\mathbf{q}^{g}))}{\sum_{R^{\prime}\in\mathcal{L}}\exp(\mathbf{u}_{R^{\prime}}^{\top}\tanh(\mathbf{W}_{sel}\mathbf{q}^{g}))},$
where $\mathbf{W}_{sel}\in\mathbb{R}^{d\times d}$ projects the state into the rule-selection space, and $\mathbf{u}_{R}\in\mathbb{R}^{d}$ is a rule-specific query vector.
4.3.3. Program Execution
The executed program state $\hat{\mathbf{q}}^{d}$ is the expected outcome of applying the rules:
(6) $\hat{\mathbf{q}}^{d}=\sum_{R\in\mathcal{L}}p(R\mid\mathbf{q}^{g})\cdot\mathcal{T}_{R}(\mathbf{q}^{g}).$
By explicitly executing these logic rules, ExeFuse bridges the ambiguity gap via logical inference, recovering implicit connections that pure similarity search would miss.
4.4. Stage 3: Granularity Calibration via Target Space Grounding
This stage addresses Challenge II (Granularity Misalignment). While logic execution ensures relevance, the resulting state $\hat{\mathbf{q}}^{d}$ may still be too abstract (e.g., generic concepts) to fit the DKG. To solve this, we perform Granularity Calibration by checking if the executed state $\hat{\mathbf{q}}^{d}$ can be successfully grounded into the Target State Space $\mathcal{Q}^{d}$.
4.4.1. Executability Score
We define the executability score $S_{exec}(f^{g})$ using kernel density estimation (KDE) over a set of DKG prototypes $F^{d}_{sub}\subset F^{d}$:
(7) $S_{exec}(f^{g})=\frac{1}{|F^{d}_{sub}|}\sum_{f^{d}_{j}\in F^{d}_{sub}}\exp\left(-\frac{\|\hat{\mathbf{q}}^{d}-\phi_{\text{dkg}}(f^{d}_{j})\|^{2}}{\tau}\right),$
where $\tau$ is a temperature hyperparameter. To ensure efficiency, $F^{d}_{sub}$ consists of $K$ cluster centroids derived from K-Means clustering on the DKG fact embeddings ($K\ll|F^{d}|$). Intuitively, $S_{exec}$ measures whether the transformed fact “lands” in a semantically valid region of the Target State Space.
4.4.2. Fusion Decision
The final fusion probability is computed by a fusion network $\Psi$ that integrates the structural context and the verification score:
@@ 4.4.2. Fusion Decision
(8) $p(\text{fuse}\mid f^{g})=\sigma\left(\Psi([\mathbf{q}^{g}\,;\,\hat{\mathbf{q}}^{d}\,;\,\tilde{S}_{exec}])\right),$
where $\tilde{S}_{exec}$ is the normalized executability score. Facts with $p>\delta$ are added to the final fused set $F_{fused}$.
@@ 4.5. Training Objective
We optimize the framework using a joint loss function on a dataset of aligned pairs $\mathcal{D}_{pos}$ (constructed via distant supervision) and negative samples.
Main Fusion Loss. We minimize the binary cross-entropy loss $\mathcal{J}_{main}$ for the fusion decision $p(\text{fuse})$:
(9) $\mathcal{J}_{main}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right].$
Rule Consistency Loss. To enforce valid semantic transformations, we minimize the regression loss on positive pairs:
(10) $\mathcal{J}_{rule}=\sum_{(f^{g},f^{d})\in\mathcal{D}_{pos}}\|\hat{\mathbf{q}}^{d}(f^{g})-\phi_{\text{dkg}}(f^{d})\|^{2}.$
The final objective is $\mathcal{J}=\mathcal{J}_{main}+\lambda\mathcal{J}_{rule}$.
4.6. Complexity Analysis
We analyze the inference complexity to demonstrate the scalability of ExeFuse. Let $|F^{g}|$ be the number of GKG facts, $d_{emb}$ be the embedding dimension, and $K$ be the number of prototypes. The complexity per fact consists of:
  • Latent Program Induction: $O(d_{emb}^{2})$ for the structure-aware neural predicate, dominated by matrix-vector multiplications.
  • Neuro-Symbolic Execution: $O(|\mathcal{L}|\cdot d_{emb})$ for attention calculation and element-wise affine transformations. Since $|\mathcal{L}|$ is small (typically <100<100), this step is highly efficient.
  • Target Space Grounding: $O(K\cdot d_{emb})$ for measuring the distance against $K$ prototypes in the Target State Space. Using clustered prototypes ($K\ll|F^{d}|$) avoids the prohibitive cost of traversing the entire DKG.
Consequently, the overall complexity is $O(|F^{g}|\cdot(d_{emb}^{2}+|\mathcal{L}|d_{emb}+Kd_{emb}))$. This scales linearly with the GKG size $|F^{g}|$, offering a significant efficiency advantage over traditional pair-wise matching methods that typically require quadratic $O(|F^{g}|\cdot|F^{d}|)$ operations.
4.7. Theoretical Motivation
To justify the design choices of ExeFuse, we analyze the expressive power of the neuro-symbolic mapping and the geometric interpretation of the grounding mechanism.
Proposition 0 (Expressive Completeness of Neuro-Symbolic Mapping).
Let $\mathcal{T}:\mathcal{Q}\to\mathcal{Q}$ be the ideal transition function mapping source facts to the target domain. The affine rule operators $\mathcal{L}$, combined with the non-linear neural selector $\mathcal{P}$, constitute a universal approximator. Specifically, the expected execution $\hat{q}^{d}=\sum_{R\in\mathcal{L}}p(R|q^{g})\cdot(w_{R}\odot q^{g}+b_{R})$ can approximate any continuous function on a compact set to arbitrary precision, given sufficient rules $|\mathcal{L}|$.
Analysis. Proposition 1 justifies the use of multiple discrete logic rules instead of a single projection matrix. While a single affine transformation is limited to linear mappings, our mechanism functions as a mixture of affine experts. The neural selector dynamically weights these experts, enabling the model to fit complex, non-linear decision boundaries required to bridge the semantic gap between GKG and DKG, ensuring that valid domain states are reachable within the model’s hypothesis space.
Proposition 0 (Granularity Alignment via Density Maximization).
Let $p_{\text{data}}(x)$ be the probability density function of valid facts in the target DKG space $\mathcal{Q}^{d}$. The executability score $S_{exec}(f^{g})$, defined via kernel density estimation over prototypes, acts as a proxy for the log-likelihood of the fused fact under the target distribution. Maximizing $S_{exec}(f^{g})$ is equivalent to minimizing the divergence between the generated fact embedding and the underlying data manifold of $G^{d}$.
Analysis. Proposition 2 clarifies that $S_{exec}$ is not merely a distance heuristic but a density-based verification mechanism. By optimizing this score, ExeFuse explicitly constrains the fused facts to reside in high-density regions of the target space (i.e., the valid semantic manifold of the domain). This theoretically guarantees that the generated facts are not just logically relevant but also conform to the specific granularity and structural patterns characteristic of the DKG.
Table 1. Dataset statistics. $F_{fused}$”: The total number of fused new facts (tuples) from GKG into DKG. “Structure. Sim.”: The average neighbor structure similarity of entities in DKG and GKG, as defined in (Jiang et al., 2024b). “#Rel. O.”: The proportion of overlapping relations in DKG and GKG.